### v1

# Unifying Attention Heads and Task Vectors via Hidden State Geometry in In-Context Learning

Abstract
The unusual properties of in-context learning (ICL) have prompted investigations into the internal mechanisms of large language models. Prior work typically focuses on either special attention heads or task vectors at specific layers, but lacks a unified framework linking these components to the evolution of hidden states across layers that ultimately produce the model’s output. In this paper, we propose such a framework for ICL in classification tasks by analyzing two geometric factors that govern performance: the separability and alignment of query hidden states. A fine-grained analysis of layer-wise dynamics reveals a striking two-stage mechanism—separability emerges in early layers, while alignment develops in later layers. Ablation studies further show that Previous Token Heads drive separability, while Induction Heads and task vectors enhance alignment. Our findings thus bridge the gap between attention heads and task vectors, offering a unified account of ICL’s underlying mechanisms.
1 Introduction
One of the most remarkable features of Large Language Models (LLMs) is their ability to respond to novel queries in user-desired mannerisms on-the-fly solely by learning from demonstrations provided in the input (as shown in Figure 1 (A))—without any additional training. This capability is known as In-context Learning (ICL) [Brown et al., 2020, Radford et al., 2019]. ICL has revolutionized natural language processing by reducing dependence on burdensome data collection and costly finetuning, enabling swift and seamless adaptation of LLMs to a variety of downstream tasks [Dong et al., 2024].
Due to its stark departure from traditional gradient-based learning paradigms, ICL has attracted significant academic interest. One line of research aims to develop a mechanistic understanding of ICL by analyzing the internal behavior of LLMs in ICL settings [Reddy, 2024]. Some studies highlight the importance of key structural components in LLMs, such as Induction Heads (IH) [Elhage et al., 2021, Olsson et al., 2022, Cho et al., 2025a], which retrieve information from demonstrations and perform copy-like operations within Transformer-based architectures. Other studies adopt a hidden-state-centric view, interpreting ICL as a process in which LLMs construct vector representations of the task from demonstrations (so-called task vectors [Hendel et al., 2023, Liu et al., 2024]) using the hidden states at an intermediate layer, and then use the vectors to steer hidden states and produce better next-token predictions. Both lines of work support their claims primarily through ablation and intervention experiments that assess the impact on model outputs.
Figure 1: (A) An example for ICL input. (B) In early layers, LLMs promote separability among the last tokens’ hidden state ($\bm{h}$) clusters w.r.t. the ground-truth labels of the queries through Previous Token Heads (PTHs). (C) In early layers or zero-shot scenarios, the direction where the hiddens are maximally separated is insufficiently aligned with the output direction (i.e., the difference vector of the label-token unembedding vectors), increasing cluster overlap after mapping and inducing higher classification errors, and (D) in later layers, Induction Heads (IHs) align these clusters towards the output direction, with the same underlying mechanism of task vectors.
However, treating model components and hidden states as separate entities in ICL analysis is both unnecessary and misleading. The two are inherently intertwined—components like attention heads transform hidden states layer by layer, and these intermediate updates ultimately determine the model’s output. As such, analyzing only the influence of attention heads or task vectors on the final prediction reduces ICL to a black-box process. A true mechanistic understanding of ICL requires tracing how these components progressively shape hidden states throughout the model, not just their end effects. A unified framework that synthesizes both perspectives and assesses their importance in the layer-wise evolution of hidden states is thus highly essential.
Therefore, in this work, we propose a framework as shown in Figure 1 for ICL in classification tasks by a key geometric insight: the classification performance of an LLM depends on whether the hidden states of queries with different ground-truth labels are (1) sufficiently separable, and (2) well aligned with the unembedding vectors of labels. In other words, hidden states should be separated along a direction that aligns with the difference between the unembedding vectors of different labels. To investigate this, we collect hidden states of queries in both ICL and zero-shot settings, and analyze them using a range of geometric measures to assess how ICL improves hidden states in terms of separability and alignment. Our approach operates on three levels: (1) We examine the layer-wise trends in these measures and compare zero-shot with ICL (and across different ICL settings) to quantify the effect of ICL on the hidden states geometry. (2) We observe how changes in measure values reflect the evolution of semantic information represented in the hidden states that ultimately informs model outputs in ICL. (3) We probe specific layers to examine the role of different types of attention heads and how they influence both the values and trajectories of these geometric measures.
Our experiments find that ICL improves classification performance primarily by enhancing alignment between query hidden states and unembedding vectors, with less impact on separability. Our results further suggest that ICL proceeds as a two-stage process: early layers primarily increase the separability of hidden states, while in middle-to-late layers, separability plateaus and alignment improve. The refinement of alignment is achieved through layer updates that filter out semantic information irrelevant to the task and preserve label-relevant content, thereby semantically refining the query hidden states. Such updates qualify ICL query hidden states as task vectors.
To build the bridge between the task vector and model components, we analyze the effect of attention heads on these geometric measures and find that (as demonstrated by Figure 1): (1) The early increase in separability is driven by a proliferation of Previous Token Heads (PTHs)—attention heads that attend to the previous token at each position—in early layers. (2) In middle-to-late layers, Induction Heads (IHs) enhance alignment by amplifying hidden state components along the unembedding directions of the labels. Thus, our findings not only explain why attention heads like PTHs and IHs are essential for ICL performance, but also clarify why induction head outputs empirically serve as effective task vectors [Todd et al., 2024, Hendel et al., 2023].
2 Related works
IH and PTH   Previous works on the mechanistic cause of ICL observe structural circuits in Transformers composed of a Previous Token Head (PTH) and an Induction Head (IH), where the PTH collects information at demonstration label tokens, and the IH copies it to query positions for prediction [Olsson et al., 2022, Singh et al., 2024, Song et al., 2025]. Their importance in classification has been observed in large-scale models as well [Cho et al., 2025a]. However, existing analyses focus on end-to-end effects and leave open how these heads influence query hidden states for classification.
Task vectors   An alternative theory proposes that LLMs summarize demonstrations into task vectors, which can be used to steer zero-shot hidden states to achieve ICL-level prediction accuracy. These vectors may be ICL query hidden states extracted at an intermediate layer [Hendel et al., 2023, Liu et al., 2024] or attention head outputs [Todd et al., 2024, Li et al., 2024]. Decoding task vectors through the unembedding matrix reveals that they encode task-related tokens and support semantic manipulations akin to “word2vec”-style algebra [Mikolov et al., 2013]. Despite intriguing empirical findings, explanations for their mechanism remain vague.
Geometry of hidden states   LLMs encode inputs in geometrically structured ways [Park et al., 2024]. For instance, they encode cities and countries by geographic relations [Gurnee and Tegmark, 2024] and numbers in rings under modular arithmetic [Liu et al., 2023]. For classification tasks, they can separate hidden states into clusters by attributes such as truth value [Marks and Tegmark, 2024] or toxicity [Lee et al., 2024]. However, the alignment of such clusters with the unembedding vectors of their labels—an equally crucial geometric factor—has received limited attention, with only a few studies noting suboptimal alignment at the output layer [Cho et al., 2025b, Kirsanov et al., 2025].
3 Theoretical framework
A classification task involves prompting an LLM with a query $x$ and ground-truth label $y$ (assumed to be a single token), e.g., $x$ = “I like this movie. Sentiment:” and $y$ = “positive”. The LLM with $L$ layers encodes $x$ into a vector $\bm{h}_{x}\in\mathbb{R}^{d}$ taken from the hidden state at the final token (“:”). Given a vocabulary ${\mathbb{V}}=\{v_{1},...,v_{|{\mathbb{V}}|}\}$ and output embedding matrix $\bm{E}\in\mathbb{R}^{|{\mathbb{V}}|\times d}$, the model predicts a distribution: $p(v_{j}|x)={\exp(\bm{E}_{v_{j}}\bm{h}_{x})}/{\sum_{j^{\prime}=1}^{|{\mathbb{V}}|}\exp(\bm{E}_{v_{j^{\prime}}}\bm{h}_{x})},$ where $\bm{E}_{v_{j}}$ is the unembedding vector of $v_{j}$. The model then outputs the token $v_{j}$ with the highest probability as the label for $x$.
In the ICL scenario, as shown in Figure 1 (A), $k$ demonstration-label pairs $(x_{1},y_{1}),...,(x_{k},y_{k})$ are prepended to form the prompt $[x_{1},y_{1},...,x_{k},y_{k},x]$, which modifies the hidden state of $x$ and results in a different distribution $p(v_{j}|x_{1},y_{1},...,x_{k},y_{k},x)$, potentially improving prediction accuracy.
Rather than a single query, we more often evaluate classification accuracy over a dataset of $n$ queries $\{x_{1},...,x_{n}\}$, with hidden states $\{\bm{h}_{1},...,\bm{h}_{n}\}$ and ground-truth labels $\{y_{1},...,y_{n}\}$. We will show that the classification accuracy depends on two geometric properties of the hidden state collection $\bm{H}=[\bm{h}_{i}]_{i=1}^{n}\in\mathbb{R}^{n\times d}$: (1) its separability, and (2) its alignment with the unembedding vectors.
As mentioned before, an LLM performs classification using the unembedding matrix $\bm{E}$, which is a multiclass classifier on the hidden states. We define its accuracy on the dataset as:
$\text{Acc}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(y_{i}=\mathop{\arg\max}_{v_{j}}\frac{\exp(\bm{E}_{v_{j}}\bm{h}_{i})}{\sum\limits_{j^{\prime}=1}^{|{\mathbb{V}}|}\exp(\bm{E}_{v_{j}^{\prime}}\bm{h}_{i})})=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(y_{i}=\mathop{\arg\max}_{v_{j}}\frac{\bm{E}_{v_{j}}\bm{h}_{i}}{\sum\limits_{j^{\prime}=1}^{|{\mathbb{V}}|}\bm{E}_{v_{j}^{\prime}}\bm{h}_{i}}),$ (1)
where the first equality states that classification accuracy equals the proportion of the labels being predicted as the most probable token and the second equality uses the monotonicity of softmax.
We first consider the simple case where $\{x_{1},...,x_{n}\}$ only has $2$ different labels, $y_{A}$ and $y_{B}$, i.e. $y_{i}\in\{y_{A},y_{B}\},\forall i$, and leave the generalization to the case of more labels to Appendix B. For a unit vector $\bm{u}\in{\mathbb{R}}^{d}$, we define the separability of $\bm{H}$ along direction $\bm{u}$ with labels $y_{A},y_{B}$ to be:
$S(\bm{u})=\frac{1}{n}(\sum_{i\in{\mathbb{N}}_{A}}\mathbbm{1}(\bm{u}^{\top}\bm{h}_{i}\geq 0)+\sum_{i\in{\mathbb{N}}_{B}}\mathbbm{1}(\bm{u}^{\top}\bm{h}_{i}<0)),$ (2)
where ${\mathbb{N}}_{A}=\{i:y_{i}=y_{A}\}$ and ${\mathbb{N}}_{B}=\{i:y_{i}=y_{B}\}$. $S(\bm{u})$ measures the fraction of hidden states that can be correctly classified by a linear decision boundary orthogonal to $\bm{u}$. By considering all possible directions, we obtain the maximum separability of $\bm{H}$, i.e. $S^{*}=\sup_{\bm{u}\in{\mathbb{R}}^{d},\|\bm{u}\|_{2}=1}S(\bm{u})$. Though $S(\bm{u})$ is not a continuous function, $S^{*}$ as a supremum can be attained on ${\mathbb{S}}^{d-1}=\{\bm{u}:\bm{u}\in{\mathbb{R}}^{d},\|\bm{u}\|_{2}=1\}$ (See Appendix A). Denote $\bm{u}^{*}$ as the $\bm{u}$ where $S$ attains the supremum, then:
Theorem 1
$\mathrm{Acc}\leq S^{*}$. The equality is achieved when $\max_{v\in{\mathbb{V}},v\notin\{y_{A},y_{B}\}}\bm{E}_{v}\bm{h}_{i}<\max(\bm{E}_{y_{A}}\bm{h}_{i},\bm{E}_{y_{B}}\bm{h}_{i}),\forall i$ and $\frac{\bm{E}_{y_{A}}-\bm{E}_{y_{B}}}{\|\bm{E}_{y_{A}}-\bm{E}_{y_{B}}\|_{2}}=c\bm{u}^{*}$ for some positive constant $c$.
This theorem shows that accuracy is upper-bounded by the maximum separability of $\bm{H}$ and that the bound is achieved when: (1) each $\bm{h}_{i}$ is maximally aligned with either $\bm{E}_{y_{A}}$ or $\bm{E}_{y_{B}}$—the unembedding vectors of $y_{A}$ and $y_{B}$—while excluding interference from other tokens, a condition we call output alignment; and (2) the direction of maximum separation $\bm{u}^{*}$ aligns with $\bm{E}_{y_{A}}-\bm{E}_{y_{B}}$, which we refer to as directional alignment. Thus, separability and alignment of hidden states critically determine accuracy. Since query hidden states in ICL differ from those in the zero-shot setting due to the presence of demonstrations, we conclude that ICL can improve classification accuracy by enhancing both their separability and alignment.
Attention heads influence query hidden states as well by their outputs to the residual stream. At layer $l$, the $h^{\text{th}}$ attention head adds $\bm{a}^{h}_{i,l}$ to the hidden state $\bm{h}_{i,l}$ of the $i^{\text{th}}$ query, affecting the transition to $\bm{h}_{i,l+1}$. Ablating a head (i.e., setting $\bm{a}^{h}_{i,l}=0$) alters all downstream hidden states and can degrade accuracy if the final-layer hidden states $\bm{H}_{L}$ lose separability or alignment.
Task-vector-based experiments operate similarly by steering hidden states at intermediate layers. In those experiments, a task vector $\bm{t}_{i}$ is added to [^1] Activation patching which directly replaces the zero-shot hidden states with ICL hidden states can also be subsumed under this form, with $\bm{t}_{i}$ being the difference between the hidden states in the two settings. $\bm{h}_{i,l}$ and the influence likewise propagates through subsequent layers. This intervention is effective if the final $\bm{H}_{L}$ has better separability and is better aligned with the labels of the task encoded in $\bm{t}_{i}$.
4 Method: Measuring separability and alignment of hidden states
Given the aforementioned influence of ICL, attention heads, and task vectors on the separability and alignment of hidden states across layers, we introduce several measures [^2] For binary classification datasets, each measure is computed over the entire $\bm{H}$. For multiclass datasets with labels $y_{1},...,y_{m}$, we enumerate all label pairs $(y_{j},y_{k})$, compute the measure using the corresponding two label clusters, i.e. hidden states of queries with labels $y_{j},y_{k}$, before averaging across all pairs to obtain a scalar summary for $\bm{H}$. As shown in Appendix B, this is justified by the fact that separability and alignment influence the accuracy of multiclass classification through the binary classification between each label pair. to quantify their influence. The formal definitions and mathematical details of all measures are provided in Appendix C.
A. Separability measure: Separability score
Theorem 1 shows that the classification is affected by the maximum separability of hidden states. However, finding maximum separability requires evaluating separability along infinitely many directions in ${\mathbb{S}}^{d-1}$ and is intractable. As a practical proxy, we train a logistic classifier on a subset of $\bm{H}$ and evaluate its accuracy on held-out data. We call the accuracy separability score as it reflects the empirical maximum separability of hidden states achieved by the classifier.
B. Alignment measures
In Theorem 1 we show that both output alignment and directional alignment affect accuracy. Below, we introduce measures for both.
B.1 Output alignment
Output alignment is defined as the classification accuracy obtained by applying the unembedding matrix $\bm{E}$ directly to hidden states $\bm{H}$, skipping subsequent layers. It measures how well $\bm{H}$ aligns with the label unembedding vectors $\bm{E}_{y_{A}}$ and $\bm{E}_{y_{B}}$, and is also known as logit lens accuracy [nostalgebraist, 2020]. When $\bm{H}$ comes from the final layer, it equals standard classification accuracy.
B.2–B.5 Directional alignment
Directional alignment is the alignment between $\bm{u}^{*}$ and the label unembedding difference direction $\frac{\bm{E}_{y_{A}}-\bm{E}_{y_{B}}}{\|\bm{E}_{y_{A}}-\bm{E}_{y_{B}}\|_{2}}$. Since finding $\bm{u}^{*}$ is intractable, we introduce the following proxies based on approximating $\bm{u}^{*}$ or measuring separability along the label-difference direction.
  • Singular alignment (B.2). To approximate $\bm{u}^{*}$, we apply Singular Value Decomposition (SVD) to the mean-centered $\bm{H}$ and extract the top-$r$ right singular vectors. These are the directions where $\bm{H}$ has the greatest spread and can approximate $\bm{u}^{*}$. We compute the maximum absolute cosine similarities between each of these vectors and $\frac{\bm{E}_{y_{A}}-\bm{E}_{y_{B}}}{\|\bm{E}_{y_{A}}-\bm{E}_{y_{B}}\|_{2}}$ as a proxy for directional alignment.
  • Variance-based alignment (B.3). To measure separability along $\frac{\bm{E}_{y_{A}}-\bm{E}_{y_{B}}}{\|\bm{E}_{y_{A}}-\bm{E}_{y_{B}}\|_{2}}$, we first measure the proportion of total variance in $\bm{H}$ that lies along this direction. A higher value indicates that this direction is a principal axis of variation—suggesting stronger directional alignment.
  • Mean-based alignment (B.4). We project the difference in means of label clusters—hidden states with labels $y_{A}$ and $y_{B}$—onto $\frac{\bm{E}_{y_{A}}-\bm{E}_{y_{B}}}{\|\bm{E}_{y_{A}}-\bm{E}_{y_{B}}\|_{2}}$ and divide it by the pooled projected variance. A higher value implies better separation margin along $\frac{\bm{E}_{y_{A}}-\bm{E}_{y_{B}}}{\|\bm{E}_{y_{A}}-\bm{E}_{y_{B}}\|_{2}}$ and better directional alignment.
  • Composite alignment (B.5). To unify the mean- and variance-based alignment, we define composite alignment as their product. It captures both the magnitude and the relative significance of separation along the label-difference direction, serving as a robust proxy for directional alignment.
C. Indirect measure: Effective dimension
As an indirect indicator of separability and alignment, we compute the effective dimension of $\bm{H}$—the number of directions along which the hidden states exhibit significant variance. A low effective dimension suggests that hidden states concentrate along a small number of directions, potentially those associated with the label unembedding vectors, and are thus easier to separate.
5 Experiments
Models   We experiment on the following 7 models: Llama2-7B, Llama2-13B, Llama2-70B [Touvron et al., 2023], Llama3-8B, Llama3-70B [Grattafiori et al., 2024], Gemma-2B, and Gemma-7B [Gemini Team et al., 2024]. Unless otherwise stated, we report the results on Llama2-70B.
Datasets   We conduct experiments on two major types of classification datasets: text classification (SUBJ [Wang and Manning, 2012], SST-2 [Socher et al., 2013], TREC [Li and Roth, 2002]) and natural language inference (SNLI [MacCartney and Manning, 2008], RTE [Dagan et al., 2005], and CB [De Marneffe et al., 2019]). We also use a generated dataset to explore the scenario where ground-truth labels of queries are not presented in the demonstrations (discussed in Subsection H.2).
ICL setting   We set the number of in-context demonstrations to be 8 unless otherwise stated. The demonstration samples are chosen randomly, except for one experiment that investigates the effect of using demonstrations procured by kNN retrieval [Liu et al., 2022]. For a detailed exposition of the implementation of models, datasets, prompt templates, etc., refer to Appendix D.
Figure 2: Comparison of trends in separability and alignment measures: ICL vs. zero-shot. Under ICL, a clear phase transition emerges: separability increases first and then alignment surges. The effective dimension first rises and then declines. This pattern is missing in the zero-shot setting. Accuracy gains from ICL over zero-shot are reflected in alignment, not separability measures.
5.1 Layer-wise trends of separability and alignment
ICL as a two-stage process   Figure 2 shows the values of six geometric measures at all layers of Llama2-70B in both zero-shot and ICL settings averaged across datasets (output alignment is deferred to Subsection 5.2 due to its strong link to the semantic information of hidden states; results for other models are in Appendix E). A key observation is that Figure 2 reveals a striking phase transition in the evolution along the layers of ICL hidden states. Phase 1 (Figure 1 (B)): In the early layers, directional alignment measures take low values, whereas separability increases rapidly. Phase 2 (Figure 1 (D)): In the subsequent layers, this trend reverses: separability plateaus, while all directional alignment measures spike simultaneously, which suggests that the label unembedding difference direction starts to capture more variance and separate label clusters more effectively. It causes the semantic information of query labels to be predominant in $\bm{H}$, which ultimately leads to final output alignment. These findings are consistent with prior work showing that task information emerges only after the initial layers [Sia et al., 2024]. The alternation between two phases is also captured by the effective dimension of $\bm{H}$, which first rises and then declines as alignment starts to improve, reflecting a gradual concentration of hidden states along label unembedding directions. In contrast, this transition is much weaker in the zero-shot setting, where effective dimension rises monotonously and the rise in alignment measures is not nearly as pronounced. Hence, this phase transition pattern characterizes ICL as a two-stage process: an initial phase enhancing separability, followed by a refinement phase that improves alignment.
ICL improves accuracy primarily through alignment   Another key observation is the small gap in separability scores between ICL and zero-shot settings—despite an 80% (Figure 2, upper middle) difference in classification accuracy. This suggests that the semantics-based separation of query hidden states is an inherent property of LLM inference, not a major result of ICL, and that alignment is the primary bottleneck in the zero-shot case. As visualized in Figure 1 (C), the direction of high separability in zero-shot hidden states is misaligned with the label-difference vector, as indicated by low singular and composite alignment scores. Along this direction, label clusters remain poorly separated (low mean-based alignment) and account for little variance (low variance-based alignment). These findings align with prior work showing that unembedding vectors often fail to reflect the main axes of separation in hidden state space [Cho et al., 2025b].
A Varying number of demonstrations
B KNN demonstration selection
C Symbolized demonstration labels
Figure 3: Layer-wise trends of separability and alignment measures in different ICL settings. (A) Phase transition is evident under demonstration numbers from 4 to 24. Accuracy improvements of increasing demonstrations are reflected by consistently improving alignment measures. (B) Changing the demonstration selection method to kNN retrieval preserves phase transition and improves accuracy through enhancing alignment. (C) Using uninformative demonstration labels hurts accuracy due to decreased alignment of hidden states, yet a similar phase transition is evident.
Given the volatility of ICL accuracy across settings with different demonstration numbers and input formats, we investigate whether the phase transition persists across settings, and whether accuracy differences can similarly be attributed to alignment properties of hidden states. We consider three settings: (1) Varying the number of demonstrations from 0 to 24 in increments of 4; (2) Changing the demonstration selection strategy to kNN retrieval where demonstrations with closest zero-shot embeddings to each query are chosen; (3) Replacing original demonstration labels with semantically uninformative symbols like @ and #. Figure 3 reports the results averaged across datasets.
Phase transition is a robust hallmark of ICL   Figure 3 shows that the phase transition pattern persists across all three settings. In Figure 3 (A), it is evident at all demonstration counts except 0. Figure 3 (B) confirms that principled demonstration selection (via kNN) preserves instead of interfering with the transition. Even in Figure 3 (C), where labels are replaced by symbols, the two-stage pattern in effective dimension remains—distinguishing it from zero-shot, where effective dimension rises monotonically—though the post-transition alignment surge is dampened.
Alignment explains accuracy differences across ICL settings   Accuracy differences between ICL settings are consistently reflected in alignment measures. In Figure 3 (A), increasing the number of demonstrations from 4 to 24 produces minimal changes in separability, while alignment measures improve in a consistent manner as the trend curves are shifted versions of one another. Effective dimension also decreases steadily, suggesting a higher concentration of hidden states along label-unembedding directions. Figure 3 (B) similarly shows that the accuracy gains from kNN-selected demonstrations correspond to stronger alignment and lower effective dimension. In Figure 3 (C), the accuracy drop from symbolic label replacement aligns with a flattened post-transition rise in alignment measures, resembling the case of zero-shot.
5.2 Semantic interpretation of alignment dynamics
Figure 4: Dynamics of alignment measures and semantics of post-transition hidden states. (A) Strong correlation between output and directional alignment; (B) Surge in alignment measures concur with encoding of label-related semantics; (C) Post-transition layers, except for the last ones, filter out unrelated semantics and retain the related ones. Refer to Appendix F for more semantics decoding cases.
To understand how the surge in directional alignment measures after the phase transition ultimately leads to high output alignment and the successful decoding of task-related labels at the final layer, we conduct a case study on SST-2 dataset to seek a semantic interpretation of the hidden states’ alignment dynamics.
Surge in alignment concurs with the encoding of label-relevant semantics   First, we compute output alignment (logit lens accuracy) across all layers and compare its trajectory to directional alignment measures. As shown in Figure 4, their strong correlation indicates that output alignment rises in sync with directional alignment. This suggests that as alignment improves, label-relevant semantics are injected into hidden states via consecutive layer updates, enabling correct label decoding even at intermediate layers. To support this, we decode the top right singular vectors of hidden states near the transition point. Figure 4 (B) reveals a sharp semantic shift: while layer 20 produces irrelevant tokens, the SST-2 label token negative emerges as the second most probable token in layer 21—immediately post-transition. This confirms that post-transition layers elevate label-relevant directions in both the semantic and variance structure of the hidden states.
Layer-wise pattern of semantic retention and filtering   To better understand how exactly layer updates after the phase transition encode relevant semantics into hidden states, we analyze the semantic information retained or removed by each layer. Given the SVDs of the centered hidden states at two consecutive layers $\bar{\bm{H}}_{l-1}=\bm{U}_{l-1}\bm{\Sigma}_{l-1}\bm{V}_{l-1}^{\top}$ and $\bar{\bm{H}}_{l}=\bm{U}_{l}\bm{\Sigma}_{l}\bm{V}_{l}^{\top}$, we compute the projection of the previous layer’s singular directions onto the next layer’s as $\bm{V}_{l-1}^{\top}\bm{V}_{l}\bm{V}_{l}^{\top}\in\mathbb{R}^{n\times d}$. The rows of this matrix with larger norms are the principal directions of $\bm{H}_{l-1}$ that are retained in $\bm{H}_{l}$, while those with smaller norms indicate filtered directions. By decoding these directions, we identify the semantic information retained or removed by each layer. The results are reported in Figure 4 (C). In most layers after the phase transition, label-related semantic directions are consistently retained and unrelated directions are filtered, leading to a progressive amplification of label-relevant information in the hidden states and a suppression of the irrelevant semantics. However, in the final layers, this pattern reverses: label-related directions are filtered, and label-irrelevant semantics are reintroduced. This shift corresponds to the slight drop in output alignment in the final layers in Figure 4 (A), and aligns with prior findings that late layers in LLMs tend to encode high-frequency, semantically uninformative tokens [Sharma et al., 2023], which may interfere with the expression of task-relevant content. The same phenomenon also occurs in other datasets as presented in Subsection F.4.
Figure 5: Accuracy gains of rank-10 denoising. See Subsection F.5 for results in other settings.
Low-rank denoising enhances label-relevant semantics   To better understand the retention-and-filtering process, we test whether filtering can be accelerated—and label-relevant semantics enhanced—by denoising the hidden states. Specifically, we apply a rank-10 SVD approximation to the centered hidden states $\bar{H}$, retaining only the top 10 directions of variation, and add the mean back. This denoising should theoretically remove less informative components and increase the prominence of label-relevant semantics. We then feed the denoised hidden states into the unembedding layer and measure output accuracy. As shown in Figure 5, low-rank denoising yields accuracy gains exceeding 10%, especially in early post-transition layers where irrelevant semantics are just beginning to be filtered. These results validate low-rank denoising and support semantic retention and filtering as a core mechanism of post-transition layers.
Task Vector properties of post-transition hidden states   As visualized by Figure 1 (D), the simultaneous enhancement of both alignment properties and label-relevant semantics after the phase transition qualifies ICL query hidden states in middle-to-late layers as task vectors [Hendel et al., 2023]. Due to their improved alignment properties, steering or replacing zero-shot hidden states with ICL hidden states can align zero-shot hidden states better with the label unembedding vectors. Similarly, modifying zero-shot hidden states with ICL hidden states, which are label-informative, makes them more suited for the prediction of the correct labels.
5.3 Significance of critical attention heads
PTHs and IHs are attention heads with distinctive attention patterns. Given a token sequence like [A][B1][A][B2]…[A][Bn][A], a PTH at the final [A] attends to the immediately preceding [Bn], while an IH attends to tokens such as [B1]–[Bn] following earlier [A] positions. For example, in the ICL prompt “I don’t like it. Answer: negative. I like it. Answer:”, at the final “:”, a PTH attends to the preceding “Answer”, while an IH attends to “negative” after the earlier “:”, effectively copying semantics of the demonstration label into the query’s hidden state. By analyzing their effects on separability and alignment across layers, we gain a more fine-grained understanding of their role in ICL—beyond simply measuring their impact on final predictions.
For each dataset and model, we rank all attention heads by their PTH and IH scores, which quantify how strongly they exhibit corresponding attention patterns on test prompts. We then ablate (zero out) the outputs of the top 10% identified IHs and PTHs and measure the resulting effects on hidden state separability and alignment across layers as well as the output accuracy. As a control, we repeat the procedure with an equal number of randomly selected heads (excluding the top PTH/IH heads). Details on identifying PTHs and IHs are provided in Appendix G.
Figure 6: Effects of attention heads ablation on the layer-wise separability and alignment measures. Ablating PTH significantly reduces separability, and ablating IH significantly reduces alignment.
PTHs induce separability, IHs induce alignment   The results in Figure 6 confirm that PTHs and IHs are crucial not only for verbalizing ICL outputs [Cho et al., 2025a] but also for structuring hidden states both separability-wise and alignment-wise. Ablating random attention heads has a negligible effect (as shown by overlapping black and gray curves). In contrast, ablating PTHs or IHs significantly alters the measures, but with distinct effects. Ablating PTHs substantially reduces separability and also influences mean-based alignment (and thus composite alignment) which concerns the separation margin of the label clusters, but has a minimal impact on other alignment measures. Conversely, ablating IHs leaves separability largely intact but severely disrupts all alignment measures, and has a far greater impact on output alignment as measured by the accuracy.
Figure 7: Average effect across datasets of steering zero-shot hidden states using task vectors created from IH outputs.
These results not only support the presence of a two-stage ICL process—initially driven by separability, then by alignment—but also clarify its mechanism. PTHs, concentrated in early layers [Cho et al., 2025a], enhance separability by attending to query tokens and encoding the distinct semantics of different queries into the query hidden states. IHs, which appear after a critical depth (30–40% of total layers [Halawi et al., 2024]), then copy demonstration label embeddings into the query hidden states through their outputs into the residual stream and improve their alignment with label unembedding directions. Moreover, in Subsection H.2 we show that PTH and IH contribute to ICL with the same mechanism even if the ground-truth label of query, which was considered a requirement for IH [Cho et al., 2025a], is not present in demonstrations.
IH outputs as task vectors   Our findings also clarify why IH outputs serve as effective task vectors, as observed by Todd et al. [2024]. Since IH outputs can improve the alignment qualities of the hidden states, they can (1) boost zero-shot accuracy when added to zero-shot hidden states, and (2) decode label-related tokens, as they align with label unembedding directions. To validate this conclusion on our classification datasets (not considered in the original study), we extract task vectors from the outputs of the identified IHs and add them to layer-30 zero-shot hidden states of Llama2-70B (see Appendix I for details). The substantial accuracy gains reported in Figure 7 support this conclusion.
6 Conclusions and Limitations
Conclusions   This work unifies two major perspectives on in-context learning (ICL)—the roles of special attention heads and task vectors—within a geometric framework centered on the separability and alignment of hidden states. We theoretically show that these two properties fully determine ICL classification accuracy and empirically demonstrate that ICL improves accuracy over zero-shot primarily by enhancing alignment, explaining why ICL query hidden states serve effectively as task vectors. Our analysis reveals a phase transition: early layers boost separability, while later layers refine alignment with label unembedding vectors by filtering out label-irrelevant semantics and preserving task-relevant directions, which coincides with the emergence of label semantics in the hidden states. By zooming into layers, we identify the complementary roles of PTHs and IHs: PTHs, prevalent in early layers, enhance separability, while IHs, emerging post-transition, promote alignment—clarifying both the phase transition and the effectiveness of IH outputs as task vectors.
Limitations   Our framework is tailored to classification and may not extend to generation or multi-token outputs. In addition, our analysis focuses on inference-time behavior in pretrained models. Future work should investigate how separability and alignment emerge during training—e.g., at what stage do models begin to separate queries by label, and when does the phase transition arise?
Appendices
Appendix A Proofs related to Theorem 1
We first show that $S^{*}$ is attained at some point $v^{*}$ in the hypersphere ${\mathbb{S}}^{d-1}=\{\bm{u}:\bm{u}\in{\mathbb{R}}^{d},\|\bm{u}\|_{2}=1\}$ though ${\mathbb{S}}^{d-1}$ is an infinite set and $S(\bm{u})$ is not continuous. Observe that each $\bm{h}_{i}$ in $\bm{H}$ partitions ${\mathbb{S}}^{d-1}$ into two hyperhemispheres. One is $\{\bm{u}:\bm{u}\in{\mathbb{R}}^{d},\|\bm{u}\|_{2}=1,\bm{u}^{T}\bm{h}_{i}\geq 0\}$ and the other is $\{\bm{u}:\bm{u}\in{\mathbb{R}}^{d},\|\bm{u}\|_{2}=1,\bm{u}^{T}\bm{h}_{i}<0\}$. Hence, as long as the number of queries in the dataset $n$ is finite, $\bm{h}_{1},...,\bm{h}_{n}$ will partition ${\mathbb{S}}^{d-1}$ into a finite number of “cells”. Each cell corresponds to the elements in ${\mathbb{S}}^{d-1}$ that satisfy a certain sign sequence $\left[\mathrm{sgn}(\bm{u}^{T}\bm{h}_{1}),...,\mathrm{sgn}(\bm{u}^{T}\bm{h}_{n})\right]\in\{1,0,-1\}^{n}$. Hence, $S(\bm{u})$ is constant on each of the cells, and the supremum (or the maximum) value $S^{*}$ is attained on one of the cells.
We then show the proof of Theorem 1:
Proof First, consider the case where ${\mathbb{V}}$ only consists of $y_{A}$ and $y_{B}$. Then $\bm{E}\in{\mathbb{R}}^{2\times d}$ and the two rows are $\bm{E}_{y_{A}}$ and $\bm{E}_{y_{B}}$. Then we have
$\text{Acc}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(y_{i}=\mathop{\arg\max}_{v_{j}}\frac{\bm{E}_{v_{j}}\bm{h}_{i}}{\sum\limits_{v_{j}^{\prime}\in\{y_{A},y_{B}\}}\bm{E}_{v_{j}^{\prime}}\bm{h}_{i}})$ (3)
$=\frac{1}{n}(\sum_{i\in{\mathbb{N}}_{1}}\mathbbm{1}(\bm{E}_{y_{A}}\bm{h}_{i}\geq\bm{E}_{y_{B}}\bm{h}_{i})+\sum_{i\in{\mathbb{N}}_{2}}\mathbbm{1}(\bm{E}_{y_{A}}\bm{h}_{i}<\bm{E}_{y_{B}}\bm{h}_{i})$ definition of argmax (4)
$=\frac{1}{n}(\sum_{i\in{\mathbb{N}}_{1}}\mathbbm{1}[(\bm{E}_{y_{A}}-\bm{E}_{y_{B}})\bm{h}_{i}\geq 0]+\sum_{i\in{\mathbb{N}}_{2}}\mathbbm{1}[(\bm{E}_{y_{A}}-\bm{E}_{y_{B}})\bm{h}_{i}<0]$ reorganizing terms (5)
$=\frac{1}{n}(\sum_{i\in{\mathbb{N}}_{1}}\mathbbm{1}[(\frac{\bm{E}_{y_{A}}-\bm{E}_{y_{B}}}{\|\bm{E}_{y_{A}}-\bm{E}_{y_{B}}\|_{2}})\bm{h}_{i}\geq 0]+\sum_{i\in{\mathbb{N}}_{2}}\mathbbm{1}[(\frac{\bm{E}_{y_{A}}-\bm{E}_{y_{B}}}{\|\bm{E}_{y_{A}}-\bm{E}_{y_{B}}\|_{2}})\bm{h}_{i}<0]$ linearity (6)
$\leq\sup_{\bm{u}\in R^{d},||\bm{u}||_{2}=1}\frac{1}{n}(\sum_{i\in{\mathbb{N}}_{1}}\mathbbm{1}(\bm{u}^{T}\bm{h}_{i}\geq 0)+\sum_{i\in{\mathbb{N}}_{2}}\mathbbm{1}(\bm{u}^{T}\bm{h}_{i}<0))$ supremum (7)
$=S^{*}.$ (8)
Here, when $\bm{E}_{y_{A}}\bm{h}_{i}=\bm{E}_{y_{B}}\bm{h}_{i}$, without loss of generality, we break the tie by letting the predicted label to be $y_{A}$ for $x_{i}$.
When there are more tokens in ${\mathbb{V}}$, it holds true that $\mathbbm{1}(y_{i}=\mathop{\arg\max}\limits_{v_{j}}\frac{\bm{E}_{v_{j}}\bm{h}_{i}}{\sum_{v_{j}^{\prime}\in\{y_{A},y_{B}\}}\bm{E}_{v_{j}^{\prime}}\bm{h}_{i}})\leq\mathbbm{1}(y_{i}=\mathop{\arg\max}\limits_{v_{j}}\frac{\bm{E}_{v_{j}}\bm{h}_{i}}{\sum\limits_{v_{j}\in 1,...,|{\mathbb{V}}|}\bm{E}_{v_{j}^{\prime}}\bm{h}_{i}})$. Hence, we have $Acc=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(y_{i}=\mathop{\arg\max}\limits_{v_{j}}\frac{\bm{E}_{v_{j}}\bm{h}_{i}}{\sum\limits_{v_{j}\in 1,...,|{\mathbb{V}}|}\bm{E}_{v_{j}^{\prime}}\bm{h}_{i}})\leq S^{*}$.
By the definition of $\bm{u}^{*}$, we have
$\frac{1}{n}(\sum_{i\in{\mathbb{N}}_{1}}\mathbbm{1}(\bm{u}^{*,T}\bm{h}_{i}\geq 0)+\sum_{i\in{\mathbb{N}}_{2}}\mathbbm{1}(\bm{u}^{*,T}\bm{h}_{i}<0))=S^{*}.$ (9)
If $\frac{\bm{E}{y_{A}}-\bm{E}{y_{B}}}{|\bm{E}{y_{A}}-\bm{E}{y_{B}}|_{2}}=c\bm{u}^{*}$ for some positive constant $c$, then
$\frac{1}{n}(\sum_{i\in{\mathbb{N}}_{1}}\mathbbm{1}[(\bm{E}_{y_{A}}-\bm{E}_{y_{B}})^{T}\bm{h}_{i}\geq 0]+\sum_{i\in{\mathbb{N}}_{2}}\mathbbm{1}[(\bm{E}_{y_{A}}-\bm{E}_{y_{B}})^{T}\bm{h}_{i}<0)])$ (10)
$=\frac{1}{n}(\sum_{i\in{\mathbb{N}}_{1}}\mathbbm{1}(\bm{u}^{*,T}\bm{h}_{i}\geq 0)+\sum_{i\in{\mathbb{N}}_{2}}\mathbbm{1}(\bm{u}^{*,T}\bm{h}_{i}<0)).$ (11)
We already showed that
$\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(y_{i}=\mathop{\arg\max}_{v_{j}}\frac{\bm{E}_{v_{j}}\bm{h}_{i}}{\sum_{v_{j}^{\prime}\in\{y_{A},y_{B}\}}\bm{E}_{v_{j}^{\prime}}\bm{h}_{i}})$ (12)
$=\frac{1}{n}(\sum_{i\in{\mathbb{N}}_{1}}\mathbbm{1}[(\bm{E}_{y_{A}}-\bm{E}_{y_{B}})^{T}\bm{h}_{i}\geq 0]+\sum_{i\in{\mathbb{N}}_{2}}\mathbbm{1}[(\bm{E}_{y_{A}}-\bm{E}_{y_{B}})^{T}\bm{h}_{i}<0)].$ (13)
When $\max_{v\in{\mathbb{V}},v\notin{y_{A},y_{B}}}\bm{E}_{v}\bm{h}_{i}<\max(\bm{E}_{y_{A}}\bm{h}_{i},\bm{E}_{y_{B}}\bm{h}_{i}),\forall i$, we have
$\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(y_{i}=\mathop{\arg\max}_{v_{j}}\frac{\bm{E}_{v_{j}}\bm{h}_{i}}{\sum\limits_{v_{j}^{\prime}\in\{y_{A},y_{B}\}}\bm{E}_{v_{j}^{\prime}}\bm{h}_{i}})$ (14)
$=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(y_{i}=\mathop{\arg\max}_{v_{j}}\frac{\bm{E}_{v_{j}}\bm{h}_{i}}{\sum\limits_{v_{j}\in 1,...,|{\mathbb{V}}|}\bm{E}_{v_{j}^{\prime}}\bm{h}_{i}})$ (15)
[…]
Appendix B Generalization of Theorem 1
In this section [^3] Here, we focus on the multiway classification, thus we change our notation to label categories from $y_{A}$ and $y_{B}$ into $y_{m}$, where $m\in\mathbb{Z}^{+}$., we consider a generalization of Theorem 1 where $x_{1},...,x_{n}$ are allowed to have $m<n$ labels, i.e. $y_{i}\in\{y_{1},...,y_{m}\},\forall i$. Same as before, let ${\mathbb{N}}_{j}=\{i:y_{i}=y_{j}\},j=1,..,m$. ${\mathbb{N}}_{j},j=1,...,m$ are pairwise disjoint and $\bigcup_{j=1}^{m}|{\mathbb{N}}_{j}|=\{1,...,n\}$. Let $S_{j,k}(\bm{u})=\frac{1}{|{\mathbb{N}}_{j}|+|{\mathbb{N}}_{k}|}(\sum_{i\in{\mathbb{N}}_{j}}\mathbbm{1}(\bm{u}^{T}\bm{h}_{i}\geq 0)+\sum_{i\in{\mathbb{N}}_{k}}\mathbbm{1}(\bm{u}^{T}\bm{h}_{i}<0))$ and let $\bm{u}^{*}_{j,k}=\mathop{\arg\max}_{\bm{u}\in{\mathbb{S}}^{d-1}}S_{j,k}(\bm{u})$ for $j,k\in 1,...,m$, and $S_{j,k}(\bm{u}_{j,k}^{*})=S^{*}_{j,k}$. For each $j$ and $k$, $\bm{u}^{*}_{j,k}$ is guaranteed to be in ${\mathbb{S}}^{d-1}$ because only $|{\mathbb{N}}_{j}|+|{\mathbb{N}}_{k}|<n$ hyperplanes are involved in separating separating ${\mathbb{S}}^{d-1}$, resulting in a smaller number of cells (guaranteed to be finite) compared to the case of $x_{1},...,x_{n}$ having only $y_{1},y_{2}$ as labels. Then, we have the following result.
Theorem 2
$\mathrm{Acc}\leq\frac{1}{n(m-1)}(\sum_{k=1}^{m}\sum_{j=k+1}^{m}(|{\mathbb{N}}_{k}|+|{\mathbb{N}}_{j}|)S^{*}_{j,k})$. The equality is achieved when for all unordered pairs $(j,k),j,k\in 1,...,m$, $\max_{v\in{\mathbb{V}},v\notin\{y_{j},y_{k}\}}\bm{E}_{v}\bm{h}_{i}<\max(\bm{E}_{y_{j}}\bm{h}_{i},\bm{E}_{y_{k}}\bm{h}_{i}),\forall i\in N_{j}\cup N_{k}$ and $\bm{E}_{y_{j}}-\bm{E}_{y_{k}}=c_{j,k}\bm{u_{j,k}}^{*}$ for some positive constant $c_{j,k}$.
The theorem states that in the case of the queries having more than two labels, the dataset classification accuracy is controlled by the pairwise separability and alignment properties of the hidden states. After calculating the separability of the hidden states of queries with labels $y_{j}$ and $y_{k}$ for each $(j,k)$, the dataset classification accuracy is then upper-bounded by these separability values weighted by the number of queries with different labels.
Similarly, the accuracy can match this upper bound if each pairwise hidden states collection is maximally separated along a direction parallel to the difference of the unembedding vectors of the corresponding labels. This theorem shows that the otherwise difficult multiclass classification problem is geometrically equivalent to the combination of a series of binary classification subproblems. And the geometric properties of the multiclass classification problem can be understood by studying the much more tractable geometry of the constituent binary subproblems.
Proof First, note the following relations
$\mathrm{Acc}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(y_{i}=\mathop{\arg\max}_{v_{j}}\frac{\bm{E}_{v_{j}}\bm{h}_{i}}{\sum\limits_{j^{\prime}\in 1,...,|{\mathbb{V}}|}\bm{E}_{v_{j}^{\prime}}\bm{h}_{i}})$
$\leq\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(y_{i}=\mathop{\arg\max}_{v_{j}}\frac{\bm{E}_{v_{j}}\bm{h}_{i}}{\sum\limits_{v_{j}^{\prime}\in\{y_{1},...,y_{k}\}}\bm{E}_{v_{j}^{\prime}}\bm{h}_{i}})$
$=\frac{1}{n}(\sum_{k=1}^{m}\sum_{x_{i}\in{\mathbb{N}}_{k}}\mathbbm{1}(\bm{E}_{y_{k}}\bm{h}_{i}\geq\bm{E}_{y_{j}}\bm{h}_{i},\forall j\neq k,j\in 1,...,m))$
$=\frac{1}{n}(\sum_{k=1}^{m}\sum_{x_{i}\in{\mathbb{N}}_{k}}\min_{j\neq k,j\in 1,...,m}\mathbbm{1}(\bm{E}_{y_{k}}\bm{h}_{i}\geq\bm{E}_{y_{j}}\bm{h}_{i}))$ indicator function is either 1 or 0
$\leq\frac{1}{n}(\sum_{k=1}^{m}\sum_{x_{i}\in{\mathbb{N}}_{k}}\frac{1}{m-1}\sum\limits_{j,j\neq k,j\in 1,...,m}\mathbbm{1}(\bm{E}_{y_{k}}\bm{h}_{i}\geq\bm{E}_{y_{j}}\bm{h}_{i}))$
$=\frac{1}{n(m-1)}(\sum_{k=1}^{m}\sum\limits_{j,j\neq k,j\in 1,...,m}\sum_{x_{i}\in{\mathbb{N}}_{k}}\mathbbm{1}(\bm{E}_{y_{k}}\bm{h}_{i}\geq\bm{E}_{y_{j}}\bm{h}_{i}))$
$=\frac{1}{n(m-1)}(\sum_{k=1}^{m}\sum\limits_{j,j\neq k,j\in 1,...,m}\sum_{x_{i}\in{\mathbb{N}}_{k}}\mathbbm{1}(\frac{\bm{E}_{y_{k}}-\bm{E}_{y_{j}}}{\|\bm{E}_{y_{k}}-\bm{E}_{y_{j}}\|_{2}}\bm{h}_{i}\geq 0)).$
Note that the sum is over all ordered pairs $(j,k)$ s.t. $j,k\in 1,...,m,j\neq k$. Consider the case of $(k=1,j=2)$ and $(k=2,j=1)$, then both $\mathbbm{1}(\bm{E}_{y_{k}}\bm{h}_{i}\geq\bm{E}_{y_{j}}\bm{h}_{i})$ and $\mathbbm{1}(\bm{E}_{y_{j}}\bm{h}_{i}\geq\bm{E}_{y_{k}}\bm{h}_{i}))$ appear in the sum. Hence, the sum can be taken over all unordered pairs $(j,k)$ and rewritten as:
$\frac{1}{n(m-1)}(\sum_{k=1}^{m-1}\sum_{j=k+1}^{m}\sum_{x_{i}\in{\mathbb{N}}_{k}\cup{\mathbb{N}}_{j}}(\mathbbm{1}(\frac{\bm{E}_{y_{k}}-\bm{E}_{y_{j}}}{\|\bm{E}_{y_{k}}-\bm{E}_{y_{j}}\|_{2}}\bm{h}_{i}\geq 0)+\mathbbm{1}(\frac{\bm{E}_{y_{k}}-\bm{E}_{y_{j}}}{\|\bm{E}_{y_{k}}-\bm{E}_{y_{j}}\|_{2}}\bm{h}_{i}<0)))$ (17)
$\leq\frac{1}{n(m-1)}(\sum_{k=1}^{m-1}\sum_{j=k+1}^{m}(|{\mathbb{N}}_{k}|+|{\mathbb{N}}_{j}|)S^{*}_{j,k}).$ (18)
Hence, we have
$\text{Acc}\leq\frac{1}{n(m-1)}(\sum_{k=1}^{m-1}\sum_{j=k+1}^{m}(|{\mathbb{N}}_{k}|+|{\mathbb{N}}_{j}|)S^{*}_{j,k}).$ (19)
As a proof of concept, plug-in $m=2$, then the bound becomes $\frac{1}{n}(|N_{1}|+|N_{2}|)S^{*}=S^{*}$ since $|N_{1}|+|N_{2}|=n$, which matches the result in Theorem 1.
To establish the equality conditions, as in the proof of Theorem 1, if $\bm{E}_{y_{j}}-\bm{E}_{y_{k}}=c_{j,k}\bm{u_{j,k}}^{*}$ for some positive constant $c_{j,k}$, then
$\sum_{i\in{\mathbb{N}}_{k}}\mathbbm{1}[(\bm{E}_{y_{k}}-\bm{E}_{y_{j}})^{T}\bm{h}_{i}\geq 0]+\sum_{i\in{\mathbb{N}}_{j}}\mathbbm{1}[(\bm{E}_{y_{k}}-\bm{E}_{y_{j}})^{T}\bm{h}_{i}<0)]$ (20)
$=\sum_{i\in{\mathbb{N}}_{k}}\mathbbm{1}(\bm{u}^{*,T}_{j,k}\bm{h}_{i}\geq 0)+\sum_{i\in{\mathbb{N}}_{j}}\mathbbm{1}(\bm{u}^{*,T}_{j,k}\bm{h}_{i}<0)=(|N_{j}|+|N_{k}|)S^{*}_{j,k}.$ (21)
We also know that if $\max_{v\in{\mathbb{V}},v\notin{y_{j},y_{k}}}\bm{E}_{v}\bm{h}_{i}<\max(\bm{E}_{y_{j}}\bm{h}_{i},\bm{E}_{y_{k}}\bm{h}_{i}),\forall i\in N_{j}\cup N_{k}$, then
$\sum_{i\in{\mathbb{N}}_{k}}\mathbbm{1}[(\bm{E}_{y_{k}}-\bm{E}_{y_{j}})^{T}\bm{h}_{i}\geq 0]+\sum_{i\in{\mathbb{N}}_{j}}\mathbbm{1}[(\bm{E}_{y_{k}}-\bm{E}_{y_{j}})^{T}\bm{h}_{i}<0)]$ (22)
$=\sum_{i\in{\mathbb{N}}_{k}\cup{\mathbb{N}}_{j}}\mathbbm{1}(y_{i}=\mathop{\arg\max}_{v_{j}}\frac{\bm{E}_{v_{j}}\bm{h}_{i}}{\sum_{v_{j}^{\prime}\in\{y_{j},y_{k}\}}\bm{E}_{v_{j}^{\prime}}\bm{h}_{i}})$ (23)
$=\sum_{i\in{\mathbb{N}}_{k}\cup{\mathbb{N}}_{j}}\mathbbm{1}(y_{i}=\mathop{\arg\max}_{v_{j}}\frac{\bm{E}_{v_{j}}\bm{h}_{i}}{\sum\limits_{v_{j}\in 1,...,|{\mathbb{V}}|}\bm{E}_{v_{j}^{\prime}}\bm{h}_{i}}).$ (24)
Finally, note that
$\text{Acc}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(y_{i}=\mathop{\arg\max}_{v_{j}}\frac{\bm{E}_{v_{j}}\bm{h}_{i}}{\sum_{v_{j}^{\prime}\in\{y_{1},y_{2}\}}\bm{E}_{v_{j}^{\prime}}\bm{h}_{i}})$ (25)
$=\frac{1}{n}\frac{1}{m-1}\sum_{k=1}^{m-1}\sum_{j=l+1}^{m}\sum_{i\in{\mathbb{N}}_{j}\cup{\mathbb{N}}_{k}}\mathbbm{1}(y_{i}=\mathop{\arg\max}_{v_{j}}\frac{\bm{E}_{v_{j}}\bm{h}_{i}}{\sum\limits_{v_{j}\in 1,...,|{\mathbb{V}}|}\bm{E}_{v_{j}^{\prime}}\bm{h}_{i}}).$ (26)
because for each ${\mathbb{N}}_{j}$, the term $\mathbbm{1}(y_{i}=\mathop{\arg\max}\limits_{v_{j}}\frac{\bm{E}_{v_{j}}\bm{h}_{i}}{\sum\limits_{v_{j}\in 1,...,|{\mathbb{V}}|}\bm{E}_{v_{j}^{\prime}}\bm{h}_{i}})$ appears $(m-1)$ times in the summation for each $i\in{\mathbb{N}}_{j}$. Hence, combining everything, we have
$\text{Acc}=\frac{1}{n}\frac{1}{m-1}\sum_{k=1}^{m-1}\sum_{j=k+1}^{m}\sum_{i\in{\mathbb{N}}_{j}\cup{\mathbb{N}}_{k}}\mathbbm{1}(y_{i}=\mathop{\arg\max}_{v_{j}}\frac{\bm{E}_{v_{j}}\bm{h}_{i}}{\sum\limits_{v_{j}\in 1,...,|{\mathbb{V}}|}\bm{E}_{v_{j}^{\prime}}\bm{h}_{i}})$ (27)
$=\frac{1}{n(m-1)}\sum_{k=1}^{m-1}\sum_{j=k+1}^{m}(|N_{j}|+|N_{k}|)S^{*}_{j,k}.$ (28)
[…]
Appendix C Calculation of geometric measures
  1. 1.
    Separability score We train the logistic classifier on half of the points in $\bm{H}$ and take its prediction accuracy on the other half as the estimated maximum separability of $\bm{H}$. We use the scikit-learn [Pedregosa et al., 2011] default implementation of the logistic classifier, and set the number of iterations to be $100$.
  2. 2.
    Output alignment We feed $\bm{H}$ into the output unembedding layer of the model to get the model’s predictions of the query labels based on $\bm{H}$ and calculate the percentage of correct predictions.
  3. 3.
    Singular alignment First obtain the centered version of $\bm{H}$, denoted as $\bar{\bm{H}}$. Then calculate its singular value decomposition, namely $\bar{\bm{H}}=\bm{U}\bm{\Sigma}\bm{V}^{T}$, and compute the absolute values of cosine similarities between the top-r singular vectors and the difference of label unembedding vectors $\bm{E}_{y_{A}}-\bm{E}_{y_{B}}$. The absolute value is taken since the sign of the singular vectors is arbitrary. In practice, we calculate the values using the top-2 singular vectors and report the maximum of the two.
  4. 4.
    Variance-based alignment First normalize $\bm{E}_{y_{A}}-\bm{E}_{y_{B}}$ as $\frac{\bm{E}_{y_{A}}-\bm{E}_{y_{B}}}{\|\bm{E}_{y_{A}}-\bm{E}_{y_{B}}\|_{2}}$. Then the ratio of variance (w.r.t. its mean) of $\bm{H}$ explained along $\frac{\bm{E}_{y_{A}}-\bm{E}_{y_{B}}}{\|\bm{E}_{y_{A}}-\bm{E}_{y_{B}}\|_{2}}$ is $\frac{\bm{E}_{y_{A}}-\bm{E}_{y_{B}}}{\|\bm{E}_{y_{A}}-\bm{E}_{y_{B}}\|_{2}}^{T}\frac{\bar{\bm{H}}^{T}\bar{\bm{H}}}{n}\frac{\bm{E}_{y_{A}}-\bm{E}_{y_{B}}}{\|\bm{E}_{y_{A}}-\bm{E}_{y_{B}}\|_{2}}$, which we take to be the variance-based alignment of $\bm{H}$.
  5. 5.
    Mean-based alignment First compute the means of label clusters (subsets of $\bm{H}$ with each of the two different labels), i.e. $\frac{1}{|{\mathbb{N}}_{A}|}\sum\limits_{i\in{\mathbb{N}}_{A}}\bm{h}_{i}$ and $\frac{1}{|{\mathbb{N}}_{B}|}\sum\limits_{i\in{\mathbb{N}}_{B}}\bm{h}_{i}$. Then take the difference and compute the value of its projection onto $\frac{\bm{E}_{y_{A}}-\bm{E}_{y_{B}}}{\|\bm{E}_{y_{A}}-\bm{E}_{y_{B}}\|_{2}}$, i.e. $(\frac{\bm{E}_{y_{A}}-\bm{E}_{y_{B}}}{\|\bm{E}_{y_{A}}-\bm{E}_{y_{B}}\|_{2}})^{T}(\frac{1}{{\mathbb{N}}_{A}}\sum\limits_{i\in|{\mathbb{N}}_{A}|}\bm{h}_{i}-\frac{1}{|{\mathbb{N}}_{B}|}\sum\limits_{i\in{\mathbb{N}}_{B}}\bm{h}_{i})$. Then, calculate the weighted average of the variances of the two label clusters along $\frac{\bm{E}_{y_{A}}-\bm{E}_{y_{B}}}{\|\bm{E}_{y_{A}}-\bm{E}_{y_{B}}\|_{2}}$, i.e. $\frac{|{\mathbb{N}}_{A}|\mathop{\mathrm{Var}}(\{\bm{h}_{i}^{T}\frac{\bm{E}_{y_{A}}-\bm{E}_{y_{B}}}{\|\bm{E}_{y_{A}}-\bm{E}_{y_{B}}\|_{2}}:i\in{\mathbb{N}}_{A}\})+|{\mathbb{N}}_{B}|\mathop{\mathrm{Var}}(\{\bm{h}_{i}^{T}\frac{\bm{E}_{y_{A}}-\bm{E}_{y_{B}}}{\|\bm{E}_{y_{A}}-\bm{E}_{y_{B}}\|_{2}}:i\in{\mathbb{N}}_{B}\})}{|{\mathbb{N}}_{A}|+|{\mathbb{N}}_{B}|}$. Finally, divide the projected difference between means by the square root of the weighted average of the projected variance.
  6. 6.
    Composite alignment Multiply the variance-based alignment by the mean-based alignment.
  7. 7.
    Effective dimension Given the SVD $\bar{\bm{H}}=\bm{U}\bm{\Sigma}\bm{V}^{T}$ and the singular values $\sigma_{1},...,\sigma_{n}$, the effective dimension of $\bm{H}$ is $\frac{{(\sum_{i=1}^{n}\sigma_{i}^{2})}^{2}}{\sum_{i=1}^{n}\sigma_{i}^{4}}$.
Appendix D Implementation details
Models   We use the official huggingface implementation of all models. All models with more than 10B parameters are quantized to 4bit.
Datasets We use the official huggingface implementation of all datasets except for the Famous People dataset crafted by ourselves. The Famous People dataset is generated using ChatGPT-4o [OpenAI et al., 2024]. The dataset has 180 datapoints, each is the name of a famous person with a profession label from one of six categories, which is one of the following six: actor, politician, singer, scientist, writer, athlete. The dataset is balanced as there are 30 data points for each label.
ICL setting   For each dataset except for the Famous People dataset, we select demonstrations from the train set and queries from the test set, or the validation set if the ground-truth labels for the test set are not provided. We keep only the first 10000 datapoints in the train set for demonstration selection if the train set has more than 10000 entries, and the first 1000 data points in the test set/validation set for accuracy evaluation. For the famous people dataset, we use the first 90 data points for demonstration selection and the remaining 90 data points for testing. For experiments involving kNN demonstration selection, we first use the respective LLMs to encode the demonstrations and queries, and select the demonstrations with embeddings closer to that of the query in $l_{2}$ distance for each query. We use the prompt formats detailed in Table 1.
Devices   All experiments with Llama2-7B, Llama2-13B, Gemma-2B, and Gemma-7B are conducted with an A800 GPU. All experiments with Llama3-8B, Llama2-70B, and Llama3-70B are conducted with an H200 GPU.
Table 1: Prompt templates and labels for different datasets.
Dataset Template Label
SST-2 {Sentence} Sentiment: {Label} positive / negative
SUBJ {Sentence} Type: {Label} subjective / objective
TREC Question: {Sentence} Type: {Label} abbreviation / entity / description / human / location / number
SNLI The question is: {Premise}? The answer is: {Hypothesis} {Label} true / maybe / false
RTE The question is: {Premise}? The answer is: {Hypothesis} {Label} true / false
CB The question is: {Premise}? The answer is: {Hypothesis} {Label} true / maybe / false
Famous People {Person Name} is a: {Label} actor / politician / singer / scientist / writer / athlete
For the label mapping experiment in Subsection 5.1. We use the mappings listed in Table 2 to map the ground-truth labels of each dataset to symbols.
Table 2: Mappings used to replace ground-truth labels of different datasets to symbols
Dataset Label Mapping
SST-2 negative/positive $\rightarrow$ @/#
SUBJ objective/subjective $\rightarrow$ @/#
TREC abbreviation / entity / description / human / location / $\rightarrow$ @/#/\$/&/*
SNLI true/maybe/false $\rightarrow$ @/#/!
RTE true/false $\rightarrow$ @/#
CB true/maybe/false $\rightarrow$ @/#/!
Appendix E Supplementary materials for Subsection 5.1
E.1 Replication of Figure 2 for other models
In Figure 8-13, we provide the visualizations of the experiments presented in Figure 2 for other models, from which similar conclusions to those in Subsection 5.1 can be drawn. Alignment measures explain the accuracy difference between ICL and zero-shot across models, and the layer-wise trends of the measures explain the phase transition pattern for all models.
E.2 Replication of Figure 3 for other models
In Figure 14-19 we provide the visualizations of the experiments presented in Figure 3 for other models to establish the generality of our conclusions. It is clear that the phase transition pattern is evident in different ICL settings across models, and alignment measures capture accuracy differences across ICL settings for all models.
Appendix F Supplementary materials for Subsection 5.2
F.1 Replication of subplot (A) of Figure 4 for other datasets and models
In Figure 2025 we provide the visualizations of the experiments presented in subplot (A) of Figure 4, to establish the generalizability of our conclusion that the directional alignment and the encoding of label-related semantics into the hidden states happen almost simultaneously after the phase transition. The layer-wise trends of the directional alignment measures and output alignment exhibit high correlation across all datasets and models.
F.2 Replication of subplot (B) of Figure 4 for other datasets
In Table 3-7, we present the tokens decoded from top singular directions near the phase transition breakpoint for other datasets. The results demonstrate that the sudden .semantic shift toward label-related tokens in the hidden states (as demonstrated by subplot (B) of Figure 4) is present in other datasets as well. Note that the phase transition happens at different layers for different datasets.
F.3 Replication of subplot (B) of Figure 4 with replaced demonstration labels
To further examine the relationship between alignment and the encoding of task-related semantics, we design a scenario in which the ground-truth demonstration labels are consistently mapped to symbols, and the expected predictions for the query label becomes the mapped symbols as well. This forces a separation between semantic encoding and alignment, as the task-related semantics remain unchanged, but alignment must now occur with respect to the unembedding vectors of the symbols. Specifically, we conduct experiments on SST-2 and SNLI where we map the SST-2 labels (negative/positive) to @ and #, and the SNLI labels (true/maybe/false) to @, #, and ! within the demonstrations. We then inspect the tokens decoded from the top singular directions of the hidden states at each layer.
As shown in Appendix J, the results reveal an interesting pattern. The first major shift in the layer-wise semantic content occurs at layer 24, where the ground-truth label tokens suddenly become dominant. This closely mirrors the pattern observed with normal demonstration labels in Table 5, and aligns with the point where output alignment begins to rise in Figure 20 (C). However, another drastic shift is observed at layer 45, where the decoded top tokens begin to include the symbol tokens. These symbols suddenly appear and start to compete with semantically relevant task tokens, marking the onset of alignment between the hidden states and the unembedding vectors of the symbols. This delayed alignment corresponds with the pattern seen in subplot (C) of Figure 3, where the phase transition under demonstrations with symbolized labels is postponed to later layers.
The competition between the true labels and the injected symbol labels persists through to the final layers. This likely contributes to the reduced accuracy in this setting, as the semantic influence of the substituted labels fails to fully override that of the ground-truth labels. Similar results for SST-2 are presented in Table 9, showing similar trends.
F.4 Replication of subplot (C) of Figure 4 for other datasets
Subplot (C) of Figure 4 reveals a two-stage pattern in the layer-wise retainment and filtering of semantic information, where most post-transition layers retain label-related semantics and filter out irrelevant ones, and the final layers do the opposite. We show in Table 10-14 that it exists in other datasets besides SST-2 as well.
F.5 Replication of Figure 5 for other datasets and models
We first provide the results of rank-10 approximations on all 6 datasets for Llama2-70B in Figure 39. The conclusion is that low-rank denoising significantly increase the output alignment significantly acorss all datasets, particularly in the innitial layers after the phase transition. This confirms the generality of our findings in Subsection 5.2. Additionally, in the final layers, the output alignment also increases on on SST-2, SUBJ, TREC, and SNLI, though not on RTE and CB.
In Figure 40-45, we present results for other models. These are consistent with the findings from Llama2-70B. Specifically, the pattern that low-rank denoising improves output alignment more in text classification datasets than in natural language inference datasets holds across models. This suggests a higher level of linguistic complexity in natural language inference tasks—likely because such tasks require reasoning over both a premise and a hypothesis, whose combined semantics are more difficult to compress and more prone to information loss during low-rank projection.
F.6 Ablation experiments for the rank hyperparameter in low-rank approximations
In this section, we explore the effect of varying the number of retained ranks in rank-$r$ SVD approximations of Llama2-70B hidden states across all datasets. Specifically, we investigate how increasingly aggressive rank reductions—setting $r=1$, $2$, and $5$—impact performance. The results, shown in Figure 46, Figure 47, and Figure 48, yield several notable findings.
For binary classification tasks such as SST-2, SUBJ, and RTE, more aggressive rank reduction sometimes leads to better output alignment, particularly in layers after the phase transition where alignment with label unembedding vectors begins to emerge. In some cases, even a rank-$1$ approximation preserves or enhances alignment-related performance. However, for three-way classification tasks like SNLI and RTE, aggressive reductions (especially at $r=1$ or $2$) can cause accuracy drops in later layers. For TREC, which involves six label classes, rank-$2$ or rank-$1$ approximations degrade accuracy across all layers.
These results align with the intuition that datasets with more labels require higher-dimensional hidden state representations to preserve the necessary separability and alignment structures among label clusters.
Appendix G Experiment details concerning the identification of PTHs and IHs
For each dataset, we use the first 50 queries to identify the set of IHs and PTHs. Denote the queries as $x_{1},...,x_{50}$ each with token length $s(x_{1}),...,s(x_{50})$.
Identification of PTHs   For each $x_{i}$, the LLM will generate an attention tensor $\bm{Attn}_{i}\in{\mathbb{R}}^{L\times N_{h}\times s(x_{i})\times s(x_{i})}$, where $L$ is the total number of layers, $N_{h}$ is the number of attention heads per layer. Hence, for the $n_{h}^{th}$ attention head at layer $l$, there is a corresponding attention matrix $\bm{Attn}(l,n_{h})_{i}\in{\mathbb{R}}^{s(x_{i})\times s(x_{i})}$, where $\bm{Attn}(l,n_{h})_{i,j,k}$ is the attention of this attention head at the $k^{th}$ token to the $j^{th}$ token in $x_{i}$. The PTH score of this attention head on $x_{i}$ is thus defined as $\sum_{k=2}^{s(x_{i})-1}\bm{Attn}(l,n_{h})_{i,k-1,k}$, which measures the total attention the attention head assigns at each token position to the previous token. Consequently, the PTH score of the $n_{h}^{th}$ attention head at layer $l$ on all the queries is $\sum_{i=1}^{50}\sum_{k=2}^{s(x_{i})-1}\bm{Attn}(l,n_{h})_{i,k-1,k}$. We calculate the PTH scores for all $(l,n_{h})$ pairs and choose the top 10% attention heads as the identified PTHs.
Identification of IHs   For each $x_{i}$, randomly select 8 demonstrations and prepare them to $x_{i}$. The resultant ICL prompt (denoted as $X_{i}$ with length $s(X_{i})$), with a slight abuse of notation, takes the format of $\langle x_{i,1}\rangle:\langle y_{i,1}\rangle,...,\langle x_{i,8}\rangle:\langle y_{i,8}\rangle,\langle x_{i}\rangle:$, where $\langle x_{i,k}\rangle$ represents the sentence part of demonstration $k$ (“I like this movie. Sentiment”) and $\langle y_{i,k}\rangle$ the label (“positive”) separated from the sentence part by a colon (:), and $\langle x_{i}\rangle$ is the sentence part for the query. At the position of the final “:”, an IH will attend to tokens after the previous “:”s, i.e. the label tokens $\langle y_{i,1}\rangle,...,\langle y_{i,8}\rangle$. Let ${\mathbb{I}}_{i}$ denote the indices of the label tokens in $X_{i}$. The IH score of the $n_{h}^{th}$ attention head at layer $l$ over the 50 queries is thus $\sum_{i=1}^{50}\sum_{k\in{\mathbb{I}}_{i}}\in\bm{Attn}(l,n_{h})_{i,k,s(X_{i})}$, i.e. we are measuring the total attention a head assigns at the final “:” position to the positions of all the label tokens summed over all 50 queries. We calculate the IH scores for all $(l,n_{h})$ pairs and choose the top 10% attention heads as the identified IHs.
Appendix H Supplementary materials for Subsection 5.3
H.1 Replication of Figure 6 for other models
In Figure 26-31, we present the results of attention heads ablations across models. The results support our conclusions regarding the significance of PTHs in enhancing the separability of the ICL hidden states and the significance of IHs in aligning them with the label unembedding vectors.
H.2 Results of attention heads ablation in the ICL setting where query label is not in context
The results for other models, reported in Figure 26-31, confirm the generality of our conclusions across model sizes and architectures. To test whether these findings hold under different ICL settings as well, we replicate the attention head ablation experiments in a special ICL scenario (“In-weight Learning”) studied by Reddy [2024] and Chan et al. [2022], where none of the in-context demonstrations share the query’s label. This setting is useful for analyzing attention heads—especially IHs, which are known for copy-like behavior—when direct label copying is not possible.
To this end, we construct a Famous People dataset following the procedures in Appendix D. Each data point is a sentence in the format “{Famous Person Name} is a:", and the label is the profession of the famous person. Once combined, they form an input sequence like, e.g., “Taylor Swift is a: Singer." that can be used as either a query or a demonstration. For each query, we sample $k=4$ demonstrations, ensuring that none share the query’s profession label. We then ablate PTHs and IHs to examine their influence on model performance. As a control, we also evaluate the measures in the setting where at least one demonstration does share the query’s label, allowing us to assess how the availability of directly copyable labels affects the impact of head ablation.
The results in Figure 32 resemble those in standard ICL settings in that ablating PTHs significantly impairs separability, while ablating IHs affects alignment. However, there are notable differences. Most importantly, the accuracy drop from ablating PTHs is greater than that from ablating IHs—opposite to the standard case, where IH ablation has a larger impact. This highlights the critical role of having the query label present in the context for IHs to function effectively. The importance of label presence is further confirmed by the substantial accuracy gain observed when the query label is forced into the context.
We provide results analogous to Figure 32 for other models in Figure 3338, which demonstrate similar results.
H.3 Results of attention heads ablation in the zero-shot setting
In this section, we investigate the effect of attention head ablation on the separability and alignment measures of zero-shot hidden states. The results averaged across datasets, as provided in Figure 49, reveal that the zero-shot setting is fundamentally different from ICL. In the zero-shot setting, ablating IHs has practically no effect on either separability or alignment measures, similar to ablating random heads, because there are no demonstrations from which to copy. On the contrary, ablating PTHs substantially compromises all separability and alignment measures. This illustrates that the contribution of PTHs to the separability of hidden states is by no means restricted to the ICL setting and highlights the significance of PTHs as key structural components of LLMs.
Appendix I Experiment details of using IH outputs to construct task vectors
For each dataset, we first construct 8-shot ICL prompts using the first 50 queries. The demonstrations used are exactly the same as the ones used to evaluate the 8-shot ICL accuracy for each dataset. Then, following the procedures described in Todd et al. [2024], we obtain the average output (across the 50 prompts) of each identified top IH at the final token position, and sum the average outputs of all IHs to obtain the task vector.
In terms of the steering experiment, we add the task vector to the hidden state of the final token of each query at Layer-30 of Llama2-70B, and let the modified hidden state flow through subsequent layers. While calculating the accuracy, we exclude the first 50 queries and only evaluate accuracy on the remaining queries in the test set to ensure an independence between the task vector and the queries involved.
Appendix J Broader impacts
Understanding the internal mechanisms of large language models is increasingly critical as these models are deployed in high-stakes applications ranging from education and healthcare to legal and governmental decision-making. This work contributes to that understanding by providing a geometric framework that explains in-context learning through the lens of hidden state dynamics, specifically separability and alignment. By identifying the role of specific architectural components—such as Previous Token Heads and Induction Heads—in shaping model behavior, our study brings interpretability to a domain often criticized for opacity.
The primary positive societal impact of this research lies in its potential to improve the transparency, controllability, and safety of LLMs. Better mechanistic understanding can inform the development of more robust models that are less reliant on spurious correlations and more capable of structured generalization. For example, insights into how semantic information is injected and refined layer by layer may support diagnostic tools that detect when a model is failing to align its internal representations with task-relevant signals. Additionally, the SVD-based techniques we propose for identifying and amplifying task-relevant components could be used to improve the efficiency or reliability of model outputs, particularly in low-resource or privacy-sensitive settings where retraining is infeasible.
However, a deeper understanding of model internals also introduces risks. The ability to precisely manipulate hidden states or attention mechanisms may enable adversarial behaviors, such as constructing prompts that selectively suppress or amplify certain outputs for political or financial gain. Moreover, techniques for filtering and steering semantic content could be misused to hide bias or simulate alignment without genuine safety improvements. As such, we encourage future work to pair interpretability research with rigorous safety and ethics evaluations, and to involve interdisciplinary expertise when applying these findings to real-world systems.
Overall, we view our contributions as a step toward more interpretable and accountable AI, but we emphasize that interpretability alone is not a guarantee of ethical deployment. Responsible application of these insights requires careful consideration of both technical and societal factors.
Figure 8: Comparison of trends in separability and alignment measures of Llama3-8B hidden states between ICL and zero-shot
Figure 9: Comparison of trends in separability and alignment measures of Llama3-70B hidden states between ICL and zero-shot
Figure 10: Comparison of trends in separability and alignment measures of Llama2-7B hidden states between ICL and zero-shot
Figure 11: Comparison of trends in separability and alignment measures of Llama2-13B hidden states between ICL and zero-shot
Figure 12: Comparison of trends in separability and alignment measures of Gemma-2B hidden states between ICL and zero-shot
Figure 13: Comparison of trends in separability and alignment measures of Gemma-7B hidden states between ICL and zero-shot
A Varying number of demonstrations
B kNN Demonstration selection
C Random symbols as demonstration labels
Figure 14: Layer-wise trends of separability and alignment measures of Llama3-8B hidden states in different ICL settings
A Varying number of demonstrations
B KNN Demonstration selection
C Random symbols as demonstration labels
Figure 15: Layer-wise trends of separability and alignment measures of Llama3-70B hidden states in different ICL settings
A Varying number of demonstrations
B KNN Demonstration selection
C Random symbols as demonstration labels
Figure 16: Layer-wise trends of separability and alignment measures of Llama2-7B hidden states in different ICL settings
A Varying number of demonstrations
B KNN Demonstration selection
C Random symbols as demonstration labels
Figure 17: Layer-wise trends of separability and alignment measures of Llama2-13B hidden states in different ICL settings
A Varying number of demonstrations
B KNN Demonstration selection
C Random symbols as demonstration labels
Figure 18: Layer-wise trends of separability and alignment measures of Gemma-2B hidden states in different ICL settings
A Varying number of demonstrations
B KNN Demonstration selection
C Random symbols as demonstration labels
Figure 19: Layer-wise trends of separability and alignment measures of Gemma-7B hidden states in different ICL settings
A SUBJ
B TREC
C SNLI
D RTE
E CB
Figure 20: Output and directional alignment of Llama2-70B ICL hidden states with label unembedding vectors on various datasets.
A SUBJ
B TREC
C SNLI
D RTE
E CB
F SST-2
Figure 21: Output and directional alignment of Llama2-7B ICL hidden states with label unembedding vectors on various datasets.
A SUBJ
B TREC
C SNLI
D RTE
E CB
F SST-2
Figure 22: Output and directional alignment of Llama3-8B ICL hidden states with label unembedding vectors on various datasets.
A SUBJ
B TREC
C SNLI
D RTE
E CB
F SST-2
Figure 23: Output and directional alignment of Llama2-13B ICL hidden states with label unembedding vectors on various datasets.
A SUBJ
B TREC
C SNLI
D RTE
E CB
F SST-2
Figure 24: Output and directional alignment of Gemma-2B ICL hidden states with label unembedding vectors on various datasets.
A SUBJ
B TREC
C SNLI
D RTE
E CB
F SST-2
Figure 25: Output and directional alignment of Llama3-70B ICL hidden states with label unembedding vectors on various datasets.
Table 3: Tokens decoded from top 10 singular directions from layers near the phase transition on SUBJ
Layer Tokens
Layer 51 opinion, humor, genre, adm, facts, plot, description, explanation, pont, humor
Layer 52 objective, humor, ., adm, dram, plot, enthus, categor, pont, subject
Layer 53 objective, humor, ., plot, fact, humor, enthus, explanation, character, humor
Table 4: Tokens decoded from top 10 singular directions from layers near the phase transition on TREC
Layer Tokens
Layer 18 Ras, Float, Physics, Heimat, Zob, rela, ijd, lbl, alphabet, uther
Layer 19 Geography, name, Physics, map, >, Ele, location, Vall, alphabet, Tests
Layer 20 definitions, histor, maps, Geography, amos, inn, map, babel, etr, letters
Table 5: Tokens decoded from top 10 singular directions of layers near the phase transition on SNLI
Layer Tokens
Layer 24 imagination, inen, exceptions, hner, igin, amerikanischer, Fich, float, opposite, orf
Layer 25 dup, unknown, similarity, Dy, Sci, ete, nost, Fich, endorf, Correct
Layer 26 erde, unknown, similarity, context, apper, architecture, unity, ambigu, alg, py
Table 6: Tokens decoded from top 10 singular directions from layers near the phase transition on RTE
Layer Tokens
Layer 21 Eva, pul, ikel, orp, Mol, enk, Dal, adrat, hor, üng
Layer 22 correct, nor, urch, Boh, VP, aqu, Dal, subscribe, currency, sign
Layer 23 correction, zero, tempo, Boh, VP, lak, FA, iva, currency, trag
Table 7: Tokens decoded from top 10 singular directions from layers near the phase transition on CB
Layer Tokens
Layer 22 TR, eign, >, qual, Joan, Icon, ebol, рит, ąż, Kontrola
Layer 23 deg, Cet, Bou, SV, ando, Simon, True, lav, iből, opposition
Layer 24 negative, schap, ftrag, court, ando, iada, rib, Martí, counter, answer
Table 8: Tokens decoded from top 10 singular directions with labels replaced on SNLI
Layer Tokens
Layer 23 Tier, aro, cum, Vector, FB, Eur, Fon, pending, Dre, Fant
Layer 24 False, aten, beck, ardo, $\nicefrac{{1}}{{3}}$, Sci, attan, Kaiser, tant, Fant
Layer 25 False, Cop, cgi, Dou, $\nicefrac{{1}}{{3}}$, Ra, hur, TF, Dre, nehm
Layer 44 false, combination, inois, multiply, scenario, group, action, group, prec, Solo
Layer 45 False, #, xspace, @, #, neutral, @, large, sentiment, False
Layer 46 False, @, contr, #, @, neutral, cheer, group, ppo, monot
Layer 78 false, False, \$., @@, @, false, \$., false, False, #
Layer 79 false, !, ., #, True, <s>, !., Hash, Street, {CJK}UTF8gbsn群
Layer 80 twe, ., X, !, :#, ##, action, !, #, Street
Table 9: Tokens decoded from top 10 singular directions with labels replaced on SST-2
Layer Tokens
Layer 24 poste, Cant, Jug, Lars, >, zak, zig, geh, eg, alus
Layer 25 rott, igi, endi, noreferrer, positive, zak, Aur, Mans, ŭ, humor
Layer 26 rott, Armen, Lau, noreferrer, negative, mee, cot, tu, f, cita
Layer 40 ipage, pra, Kle, estaven, Maj, CURLOPT, extreme, humor, $\odot$, anel
Layer 41 ViewById, pra, meno, Einzelnach, ziel, @, disappoint, Meyer, L, humor
Layer 42 negative, pra, nar, Einzelnach, #, humor, extreme, humor, embargo, #
Layer 78 negative, пози, elt, neutral, hyper, \$., <s>, positive, inea, intros
Layer 79 positive, ( {, *., neutral, Tru, ???, <s>, ?, humor, Sent
Layer 80 :(, positive, $\cdot$, !, disappoint, positive, about, positive, wod, intros
Table 10: Tokens decoded from top retained and filtered directions in layer updates on SUBJ
Retained Tokens
Stage Layer Tokens
Stage 1 Layer 54 humor, description, explanation, objective, humor, shock, subject, character, comparison, humor
Layer 59 humor, plot, pra, facts, objective, hypoth, subject, met, explanation, subject
Layer 66 ., objective, Plot, explan, prom, subject, character, fact, objective, sci
Stage 2 Layer 80 plot, observation, IM, Thom, verb, hyper, quar, Em, ps, Guillaume
Filtered Tokens
Stage Layer Tokens
Stage 1 Layer 54 oro, Kant, Harr, Werner, Bast, desert, Kent, Uns, ola, Fall
Layer 59 iter, Arg, halten, enser, Lok, Kontrola, tring, atif, ias, Sof
Layer 66 conj, dens, shal, amo, Lisa, lish, Sob, narrow, UMN, co
Stage 2 Layer 80 subject, Met, chron, intros, mix, emot, humor, Type, jective, subject
Table 11: Tokens decoded from top retained and filtered directions in layer updates on TREC
Retained Tokens
Stage Layer Tokens
Stage 1 Layer 30 pione, taste, definition, temporal, precision, science, amen, lich, irectory, enta
Layer 50 description, subst, person, numbers, interpret, location, wars, location, answer, date
Layer 70 нау, date, number, location, abbre, currency, location, number, description, Creative
Stage 2 Layer 80 hol, Heidel, Desp, eth, olis, description, origin, Type, description, arto
Filtered Tokens
Stage Layer Tokens
Stage 1 Layer 30 ako, dn, amen, ainer, overlay, aar, CHAP, storm, synchron, tober
Layer 50 Gordon, Chinese, acci, Grace, penas, zak, gas, Kü, ps, prototype
Layer 70 mas, fit, ioni, Salt, ru, oci, âtre, quet, thro, Sever
Stage 2 Layer 80 , number, list, discipline, organ, entertain, weather, bum, poll, coff
Table 12: Tokens decoded from top retained and filtered directions in layer updates on SNLI
Retained Tokens
Stage Layer Tokens
Stage 1 Layer 25 pac, ício, charm, beck, hist, rap, iella, fact, contradiction, engol
Layer 50 ., probable, false, uncertain, solo, neutral, tender, unknown, gender, unlikely
Layer 70 ., reverse, false, scenario, maybe, reverse, identity, gender, conce, Gram
Stage 2 Layer 80 rola, Dic, Kra, Column, Common, motor, …, ge, Salt, kern
Filtered Tokens
Stage Layer Tokens
Stage 1 Layer 25 brief, uclide, Ath, bul, références, EC, asta, Rot, nia, Taml
Layer 50 aho, oba, zu, umble, prepared, Mey, jer, Bras, irre, Maj
Layer 70 Yu, Mare, iner, artificial, overlap, ilder, aret, elev, constru, Storm
Stage 2 Layer 80 im, <s>, fl, maybe , similarity, sem, habit, Ta, gender, action
Table 13: Tokens decoded from top retained and filtered directions in layer updates on RTE
Retained Tokens
Stage Layer Tokens
Stage 1 Layer 33 Wikipedia, disput, yes, Jakob, alias, Ven, essel, indirect, False, isu
Layer 50 false, alarm, irrelevant, opposite, nost, naming, duplicates, misunder, Geography, location
Layer 68 :(, Geography, lament, sto, impossible, opinion, ken, numer, negative, true
Stage 2 Layer 80 Ale, Ad, Hub, Ke, lub, Terra, Haz, Ta, part, Andre
Filtered Tokens
Stage Layer Tokens
[…]
Layer 50 ловi, CA, Um, wob, prot, jär, {CJK}UTF8gbsn望, ang, {CJK}UTF8gbsn孝, Roman
Layer 68 fert, CA, bach, Word, RL, gon, Charlie, MS, nen, Feld
Stage 2 Layer 80 SR, techni, gr, unclear, trag, opposite, Susan, estimate, rare, murder
Table 14: Tokens decoded from top retained and filtered directions in layer updates on CB
Retained Tokens
Stage Layer Tokens
Stage 1 Layer 25 Bou, True, <s>, Cet, ando, elin, лом, Pear, stabil, lav
Layer 49 true, unlikely, uncertain, absence, False, predictions, avoided, myth, predictions, asym
Layer 70 true, false, agreement, maybe, predictions, silent, negative, disapp, áv, predictions
Stage 2 Layer 80 Giov, future, often, Frank, tend, imet, Sid, fran, Intent, maybe
Filtered Tokens
Stage Layer Tokens
Stage 1 Layer 25 slant, acci, Laurent, cht, brie, chamber, olare, gart, una, vie
Layer 49 Dra, cito, nick, Verm, moz, neglect, abel, WR, anger, uchar
Layer 70 ottom, nation, Cooper, exhaust, reign, hem, Um, diag, Ri, wa
Stage 2 Layer 80 false, esi, , False, Yes, often, aer, “, Mod, George
Figure 26: Effects of attention heads ablation on the layer-wise separability and alignment measures of Llama3-8B hidden states averaged across all datasets
Figure 27: Effects of attention heads ablation on the layer-wise separability and alignment measures of Llama3-70B hidden states averaged across all datasets
Figure 28: Effects of attention heads ablation on the layer-wise separability and alignment measures of Llama2-7B hidden states averaged across all datasets
Figure 29: Effects of attention heads ablation on the layer-wise separability and alignment measures of Llama2-13B hidden states averaged across all datasets
Figure 30: Effects of attention heads ablation on the layer-wise separability and alignment measures of Gemma-2B hidden states averaged across all datasets
Figure 31: Effects of attention heads ablation on the layer-wise separability and alignment measures of Gemma-7B hidden states averaged across all datasets
Figure 32: Effects of attention heads ablation on the layer-wise separability and alignment measures of Llama2-70B hidden states on Famous People dataset.
Figure 33: Effects of attention heads ablation on the layer-wise separability and alignment measures of Llama3-8B hidden states on Famous People dataset.
Figure 34: Effects of attention heads ablation on the layer-wise separability and alignment measures of Llama3-70B hidden states on Famous People dataset.
Figure 35: Effects of attention heads ablation on the layer-wise separability and alignment measures of Llama2-7B hidden states on Famous People dataset.
Figure 36: Effects of attention heads ablation on the layer-wise separability and alignment measures of Llama2-13B hidden states on Famous People dataset.
Figure 37: Effects of attention heads ablation on the layer-wise separability and alignment measures of Gemma-2B hidden states on Famous People dataset.
Figure 38: Effects of attention heads ablation on the layer-wise separability and alignment measures of Gemma-7B hidden states on Famous People dataset.
A SST-2
B SUBJ
C TREC
D SNLI
E RTE
F CB
Figure 39: Accuracy gains of rank-10 denoising over layers on all datasets with Llama2-70B
A SST-2
B SUBJ
C TREC
D SNLI
E RTE
F CB
Figure 40: Accuracy gains of rank-10 denoising over layers on all datasets with Llama3-8B
A SST-2
B SUBJ
C TREC
D SNLI
E RTE
F CB
Figure 41: Accuracy gains of rank-10 denoising over layers on all datasets with Llama3-70B
A SST-2
B SUBJ
C TREC
D SNLI
E RTE
F CB
Figure 42: Accuracy gains of rank-10 denoising over layers on all datasets with Llama2-7B
A SST-2
B SUBJ
C TREC
D SNLI
E RTE
F CB
Figure 43: Accuracy gains of rank-10 denoising over layers on all datasets with Llama2-13B
A SST-2
B SUBJ
C TREC
D SNLI
E RTE
F CB
Figure 44: Accuracy gains of rank-10 denoising over layers on all datasets with Gemma-2B
A SST-2
B SUBJ
C TREC
D SNLI
E RTE
F CB
Figure 45: Accuracy gains of rank-10 denoising over layers on all datasets with Gemma-7B
A SST-2
B SUBJ
C TREC
D SNLI
E RTE
F CB
Figure 46: Accuracy gains of rank-5 denoising over layers on all datasets with Llama2-70B
A SST-2
B SUBJ
C TREC
D SNLI
E RTE
F CB
Figure 47: Accuracy gains of rank-2 denoising over layers on all datasets with Llama2-70B
A SST-2
B SUBJ
C TREC
D SNLI
E RTE
F CB
Figure 48: Accuracy gains of rank-1 denoising over layers on all datasets with Llama2-70B
Figure 49: Average effects across datasets of attention heads ablation on the layer-wise separability and alignment measures of Llama2-70B hidden states in the zero-shot setting.